# GraphXAIN: Narratives to Explain Graph Neural Networks

### v2

@@ Abstract
Graph Neural Networks (GNNs) are a powerful technique for machine learning on graph-structured data, yet they pose interpretability challenges, especially for non-expert users. Existing GNN explanation methods often yield technical outputs such as subgraphs and feature importance scores, which are not easily understood. Building on recent insights from social science and other Explainable AI (XAI) methods, we propose GraphXAIN, a natural language narrative that explains individual predictions made by GNNs. We present a model-agnostic and explainer-agnostic XAI approach that complements graph explainers by generating GraphXAINs, using Large Language Models (LLMs) and integrating graph data, individual predictions from GNNs, explanatory subgraphs, and feature importances. We define XAI Narratives and XAI Descriptions, highlighting their distinctions and emphasizing the importance of narrative principles in effective explanations. By incorporating natural language narratives, our approach supports graph practitioners and non-expert users, aligning with social science research on explainability and enhancing user understanding and trust in complex GNN models. We demonstrate GraphXAIN’s capabilities on a real-world graph dataset, illustrating how its generated narratives can aid understanding compared to traditional graph explainer outputs or other descriptive explanation methods.
@@ 1 Introduction
The exponential growth in the complexity of machine learning models has led to architectures reaching billions of parameters, resulting in significant improvements in their performance. [19, 29, 8, 3]. As these complex ‘black-box’ models, characterized by their high accuracy yet lack of interpretability, continue to evolve, the demand for transparency has intensified [26]. Explainable Artificial Intelligence (XAI) has emerged to address this challenge by enhancing the trustworthiness and transparency of complex decision-making processes [12].
@@
Graph Neural Networks (GNNs) [30] have recently gained notable success and have become state-of-the-art solutions for modelling relational data characterized by nodes connected via edges [33, 35, 15]. However, the need for interpretability in GNNs remains [36, 18, 16].
Existing GNN explanation methods primarily provide explanatory subgraphs and feature importance values, as illustrated in the upper section of Figure 1, which displays an explanation generated by GNNExplainer [36] to explain the GNN model’s prediction for node 57 in the NBA dataset, based on player’s attributes, fields statistics and social connections. Interpreting the subgraph visualization alongside feature importance metrics alone can pose significant challenges for readers. Without a complementary narrative, end users must rely solely on the technical outputs of subgraphs and feature importance to interpret the model’s underlying reasoning, which can be particularly challenging given the complex nature of GNN models. Despite its limitations, GNNExplainer remains a state-of-the-art graph explainer.
Such an approach is not aligned with the comprehensibility postulate introduced by Michalski (1983) [23],
@@
which advocates that computer-generated results should be directly interpretable in natural language. Moreover, effective explanations should enhance the alignment between the user’s mental model and AI model [21, 14]. GNNExplainer’s approach, among other graph explainers, does not adhere to these principles and exemplifies the phenomenon termed the inmates running the asylum [25], where solutions are technical and cater primarily to experts, neglecting the needs of non-technical users who prefer explanations in natural language [4]. Martens et al. (2023) [20] proposed XAIstories, a framework that employs Large Language Models (LLMs) to generate narratives explaining AI model’s predictions based on SHAP [19] and counterfactual (CF) [22, 34] explanations. In their survey among data scientists and lay users, over 90% of the general audience found these narratives convincing, while 83% of the surveyed data scientists indicated they would likely use XAIstories to communicate explanations to non-expert audiences. By integrating narrative communication as a compliment to the XAI explanations, we can present the model’s decision-making process in a manner that aligns with human cognitive preferences, enhancing explanations beyond mere descriptions and effectively illustrating cause-and-effect relationships [4].
@@
To the best of our knowledge, no previous research proposes the general framework for natural language XAI Narratives for Graph Neural Network models. In this article, we introduce the first method to automatically generate GraphXAINs, natural language narratives, for explaining GNN models. By complementing explanatory subgraphs and feature importances with coherent XAI Narratives, we aim to present the decision-making processes of GNNs in a more transparent and accessible way to end users. We posit that this framework will not only enhance interpretability but also facilitate more effective communication of model insights across various graph applications.
@@
We illustrate our approach by integrating GraphXAIN with the existing graph XAI framework (GNNExplainer) for GNNs and demonstrating its explanatory abilities on a real-world dataset.
@@ 2.1 Explainability in Machine Learning
Several approaches have been proposed to enhance the explainability of machine learning models across various modalities, including image data [29, 31, 32, 1, 11], tabular data [29, 19, 2], natural language [32, 19, 29], and unstructured data such as graphs [36, 18]. One of the most popular XAI methodologies are post-hoc explanations, which aim to interpret deep learning models after the training stage [26]. These methods encompass feature importance measures, visualisation techniques, and surrogate models. For instance, SHAP (SHapley Additive exPlanations) [19] and LIME (Local Interpretable Model-agnostic Explanations) [29] estimate the contribution of each feature to a particular prediction, thereby shedding light on the model’s decision-making process for specific instances.
@@
However, despite these advancements, challenges persist in ensuring that explanations are both methodologically accurate [7] and meaningful to end users [20, 27]. Explanations must bridge the gap between technical complexity and user comprehension, necessitating a careful balance between fidelity and interpretability [4, 25]. The need for explainability methods that are understandable to non-technical users is primarily crucial in sensitive domains such as healthcare, finance, and legal systems, where understanding the deep learning model’s decision-making process is essential for trust and transparency [29, 26].
2.2 Graph Neural Networks Explainability
@@
GNNs are increasingly used for modelling relational data in domains such as social networks, molecular structures, and knowledge graphs [30, 15, 33]. Their complex architectures, however, pose challenges for understanding and interpreting their predictions. The first method developed to address this issue is GNNExplainer [36], which provides instance-level explanations by identifying a compact subgraph and relevant node features that are most influential for a specific prediction. GNNExplainer formulates the explanation task as an optimisation problem, maximising the mutual information between the explanatory subgraph with a subset of node features, and the input graph that is subject to explain.
@@
Among other graph explanation methods, Lucic et al. (2021) [18] introduced CF-GNNExplainer which alters the GNNExplainer to answer ’what-if’ questions using the counterfactual explanation approach. Rather than merely identifying influential features or subgraphs, CF-GNNExplainer searches for minimal perturbations to the input graph that would alter the GNN model’s prediction by edge deletion. This method demonstrates how small changes in graph structure or node features impact outcomes, enhancing understanding of the model’s decision-making process.
@@
Although the aforementioned state-of-the-art graph explanation frameworks are methodologically sound, they merely provide users with additional graphs and node feature importance values that are not easily interpretable by end-users, thereby limiting their practical utility and compelling users to construct the explanatory narrative themselves.
@@
Incorporating natural language into GNN explanations could bridge the gap between technical outputs and human understanding by translating complex model reasoning into accessible narratives, thereby enhancing user comprehension and trust. However, previous methods to generate them are not tailored to the popular GNNExplainer outputs and/or provide descriptions rather than narratives (see below for a discussion of the important differences). Giorgi et al. (2024) [6] addressed this issue by using LLMs to generate textual explanations for counterfactual graphs. However, their explanations lack contextual information and do not illustrate cause-and-effect relationships, resulting in descriptive communication rather than narrative explanations [4] (see Appendix B.2 for examples).
@@
He et al. (2024) [10] used LLMs to generate natural language explanations for counterfactual graphs in the context of molecular property prediction. However, since their framework adheres to domain-specific knowledge, it cannot be considered a general method for graph model explanations. Furthermore, the explanations produced exhibit the same limitation as in Giorgi et al. (2024) [6], being more descriptive rather than narrative (see Appendix B.3 for examples).
@@
The most comparable approach to ours is presented by Pan et al. (2024) [27], who developed TAGExplainer, a method for generating natural language explanations for Text-Attributed Graphs (TAGs). Although their explanations incorporate elements of narrative communication, their framework is limited only to TAGs. This limitation restricts the generalisability of TAGExplainer to broader graph data modelling contexts, as TAGs represent only a subset of real-world graph data (see Appendix B.4 for examples). Moreover, TAGExplainer does not provide an explanatory subgraph or feature importance values.
@@ 3 Methods
3.1 XAI Narratives and Descriptions
Research in communication theory and psychology indicates that narrative-based explanations are more accessible and memorable than descriptive methods, making them effective for conveying scientific evidence to non-expert audiences [4]. Moreover, narratives are processed more rapidly by individuals without prior knowledge and are more engaging and persuasive, thereby enhancing trust and understanding of AI models [4, 25, 20].
Further, as narrative communication relies on contextual cause-and-effect relationships, it is considerably more challenging to fragment a narrative into smaller, meaningful segments without either significantly altering the interpretation of these segments or disrupting the coherence of the original narrative [4]. Consequently, narratives are often perceived as storytelling, characterised by a coherent structure comprising an introduction, main body, and conclusion. In contrast, descriptive, context-free communication can be readily fragmented into smaller units while still effectively conveying the necessary information [4]. The XAIstories [20] framework addresses aforementioned explanation limitations by enhancing the narrative communication of SHAP and CF explanations, aligning with the research on human-AI interactions [4, 25].
Having identified the need to distinguish between narrative and descriptive explanations in the context of XAI, we propose definitions for both terms, drawing on research from the social sciences [4, 25].
@@ 3.1 XAI Narratives and Descriptions
An example description would list the most important features and neighbouring nodes. This relates closely to the data-to-text or graph-to-text [28, 5]. Figure 5 provides an XAI Description (XAID) example for the XAI solution provided by GNNExplainer shown in Figure 1. Clearly, an XAID is less valuable than an XAIN, as descriptions are less accessible and memorable than narrative communication methods, making them less effective for conveying scientific evidence to a broader audience [4, 20].
3.2 From GNN models to Natural Language Narratives
To overcome the issue of purely technical explanations of GNN models’ predictions, we propose GraphXAINs—natural language narratives for graphs. We propose the following definition of GraphXAIN, a XAI Narratives for graphs:
@@ 3.2 From GNN models to Natural Language Narratives
Our solution involves converting subgraph structures and feature importance metrics derived from graph explainers into GraphXAINs - coherent natural language narratives. The workflow, presented in Figure 2, begins with the training of a GNN classification model on a relational graph dataset. Our approach is not limited to pure graphs; data can include associated node and edge features. Further, using the graph explainer method, we extract relevant subgraphs and feature importances. These technical outputs are then transformed and prompted into an LLM for further inference. The LLM generates a natural language narrative that explains the prediction of the target node’s label, highlighting the contributions of specific neighbouring nodes, corresponding edges, and node features in a context-dependent manner, resulting in a coherent cause-and-effect narrative presented to the end user.
Importantly, our framework is agnostic to the graph data type, graph model, and graph explainer, allowing its application across various graph scenarios. By translating technical explanations into coherent narratives, we aim to make graph model explanations more intuitive and accessible for practitioners.
@@ 4 Experiments
4.1 Dataset
We conducted our experiments on a real-world graph NBA dataset111https://www.kaggle.com/noahgift/social-power-nba, which includes players’ performance statistics from the NAB 2016-2017 season, alongside various personal attributes such as height, weight, age, and nationality. The graph was constructed by linking NBA players based on their relationships on the Twitter platform. After preprocessing, filtering out disconnected nodes, and ensuring the graph is undirected, where any connection on the Twitter platform results in a mutual edge, the final graph comprised of the total number of 400 nodes (players), 42 node features, and 21,242 edges (Twitter connections).
The node classification task involves predicting whether a player’s salary exceeds the median salary. Prior to training, the data were randomly divided into training, validation, and test sets in a 60/20/20 split.
4.2 Graph Model
We trained a two-layer Graph Convolutional Network (GCN) [15] model with 16 hidden channels. For training we used the Binary Cross-Entropy loss function and the AdamW [17] optimizer with a learning rate of 0.001 and a weight decay of 510-4. The training process continued through 1400 epochs, resulting in the GCN model achieving a 0.80 test AUC.

### v3

@@ Abstract
Graph Neural Networks (GNNs) are a powerful technique for machine learning on graph-structured data, yet they pose challenges in interpretability. Existing GNN explanation methods usually yield technical outputs, such as subgraphs and feature importance scores, that are difficult for non-data scientists to understand and thereby violate the purpose of explanations. Motivated by recent Explainable AI (XAI) research, we propose GraphXAIN, a method that generates natural language narratives explaining GNN predictions. GraphXAIN is a model- and explainer-agnostic method that uses Large Language Models (LLMs) to translate explanatory subgraphs and feature importance scores into coherent, story-like explanations of GNN decision-making processes. Evaluations on real-world datasets demonstrate GraphXAIN’s ability to improve graph explanations. A survey of machine learning researchers and practitioners reveals that GraphXAIN enhances four explainability dimensions: understandability, satisfaction, convincingness, and suitability for communicating model predictions. When combined with another graph explainer method, GraphXAIN further improves trustworthiness, insightfulness, confidence, and usability. Notably, 95% of participants found GraphXAIN to be a valuable addition to the GNN explanation method. By incorporating natural language narratives, our approach serves both graph practitioners and non-expert users by providing clearer and more effective explanations.
@@ 1 Introduction
The exponential growth in the complexity of machine learning models has led to architectures reaching billions of parameters, resulting in significant improvements in their performance [25, 37, 12, 5]. As these complex ‘black-box’ models, characterized by their high accuracy yet lack of interpretability, continue to evolve, the demand for transparency has intensified [33]. Explainable Artificial Intelligence (XAI) has emerged to address this challenge by enhancing the trustworthiness and transparency of complex decision-making processes [17].
@@
Graph Neural Networks (GNNs) have recently gained notable success and become state-of-the-art solutions for modelling relational data characterized by nodes connected via edges [38, 41, 43, 21]. However, the need for interpretability in GNNs remains [44, 24, 22].
Existing GNN explanation techniques predominantly offer explanatory subgraphs and feature-importance attributions, as shown in the upper section of Figure 1. The figure illustrates an explanation produced by GNNExplainer for the GNN model’s classification of node 57 in the NBA dataset, taking into account the player’s attributes, field statistics, and social connections [44]. However, the correct interpretation of the subgraph visualization alongside feature importance metrics alone can pose significant challenges for readers. Without a complementary narrative, practitioners employing this explanation technique to interpret the GNN’s predictions must rely solely on subgraphs and feature importance outputs, which can be particularly challenging given the complex nature of GNN models. Despite its limitations, GNNExplainer remains a state-of-the-art graph explainer. Such an approach is not aligned with the comprehensibility postulate introduced by Michalski (1983), which advocates that computer-generated results should be directly interpretable in natural language [28]. Moreover, effective explanations should enhance the alignment between the user’s mental model and AI model [27, 20]. GNNExplainer’s approach, among other graph explainers, does not adhere to these principles and exemplifies the phenomenon termed the “Inmates Running the Asylum”, where solutions are technical and cater primarily to experts, overlooking the needs of less-technical practitioners who prefer natural-language explanations [31, 6].
@@
Martens et al. (2023) proposed XAIstories, a framework that employs Large Language Models (LLMs) to generate narratives explaining the AI model’s predictions based on SHAP and counterfactual (CF) explanations [26]. In their survey among data scientists and lay users, over 90% of the general audience found these narratives convincing, while 83% of the surveyed data scientists indicated they would likely use XAIstories to communicate explanations to non-expert audiences. By incorporating narrative communication alongside technical XAI methods, the model’s decision-making process can be presented in a manner that aligns with human cognitive preferences, elevating explanations beyond mere descriptions and clarifying cause-and-effect relationships [6].
@@
To the best of our knowledge, no previous research proposes the general framework for natural language XAI Narratives to explain Graph Neural Network models’ predictions. In this article, we introduce the first method to automatically generate GraphXAIN, a natural language narrative, to explain the GNN model. By complementing explanatory subgraphs and feature importances with coherent XAI Narratives (see Figure 2), we aim to further explain the decision-making process of GNNs in a more transparent and accessible way. We posit that this method will not only enhance interpretability but also facilitate more effective communication of model predictions across various graph applications.
@@
We illustrate our approach by integrating GraphXAIN with the existing graph XAI framework (GNNExplainer) for GNNs and demonstrating its explanatory abilities on real-world datasets, both in classification and regression tasks.
@@ 2.1 Explainability in Machine Learning
Several approaches have been proposed to enhance the explainability of machine learning models across various modalities, including image data[37, 39, 40, 1, 15], tabular data[37, 25, 4], natural language[40, 25, 37], and unstructured data such as graphs[44, 24]. Among the most popular XAI methodologies are post-hoc explanations, which aim to interpret prediction models after the training stage [33], while counterfactual explanations indicate what minimal changes in the input data are required to obtain a different predicted class [27, 42]. These methods encompass feature importance measures, visualisation techniques, and surrogate models. For instance, SHAP (SHapley Additive exPlanations[25]) and LIME (Local Interpretable Model-agnostic Explanations[37]) estimate the contribution of each feature to a particular prediction.
@@
However, despite these advancements, challenges persist in ensuring that explanations are both methodologically accurate and meaningful for a variety of stakeholders, not only for the data scientists [10, 26, 34]. Explanations must bridge the gap between technical complexity and user comprehension, necessitating a careful balance between fidelity and interpretability[6, 31]. The need for explainability methods that are understandable to less-technical users is primarily crucial in sensitive domains such as healthcare, finance, and legal systems, where understanding the model’s decision-making process is essential for trust and transparency[37, 33].
2.2 Explainability in Graph Neural Networks
@@
GNNs are increasingly used for modelling relational data in domains such as social networks, molecular structures, and knowledge graphs [38, 21, 41]. Their complex architectures, however, pose challenges for understanding and interpreting their predictions. The first method developed to address this issue is GNNExplainer [44], which provides instance-level explanations by identifying an explanatory subgraph and relevant node features which are most influential for a specific prediction. GNNExplainer formulates the explanation task as an optimisation problem, maximising the mutual information between the explanatory subgraph with a subset of node features and the original graph that is subject to explain.
@@
Among other graph explanation methods, Lucic et al. (2021) introduced CF-GNNExplainer[24], which alters the GNNExplainer to answer ’what-if’ questions using the counterfactual explanation approach. Rather than merely identifying influential features or subgraphs, CF-GNNExplainer searches for minimal perturbations to the original graph that would alter the GNN model’s prediction by edge deletion. This method demonstrates how small changes in graph structure impact outcomes, enhancing understanding of the model’s decision-making process.
@@
Although the aforementioned state-of-the-art graph explanation frameworks are methodologically sound, they merely provide users with additional graphs and node feature importance values that are not easily interpretable (see Figure 1), thereby limiting their practical utility and compelling practitioners to construct the explanatory narrative themselves.
@@
We posit that incorporating natural language into GNN explanations could bridge the gap between technical outputs and human understanding by translating complex model reasoning into accessible narratives, thereby enhancing comprehension and trust among practitioners. However, previous methods to generate them are not tailored to the popular GNNExplainer output and/or provide descriptive explanation rather than a cohesive narrative (see below for a discussion of the important differences). Giorgi et al. (2024) addressed this issue by using LLMs to generate textual explanations for counterfactual graphs[9]. However, their explanations lack contextual information and do not illustrate cause-and-effect relationships, resulting in primarily descriptive communication rather than narrative explanation - the latter being more valuable for conveying more complex information[6] (see Appendix B.2 for examples).
@@
He et al. (2024) used LLMs to generate natural language explanations for counterfactual graphs in the context of molecular property prediction[14]. However, since their framework adheres to domain-specific knowledge, it cannot be considered a general method for graph model explanations with natural language explanations. Furthermore, the explanations produced by their method exhibit the same limitation as in Giorgi et al. (2024), being more descriptive rather than narrative (see Appendix B.3 for examples).
@@
The most comparable approach to ours is presented by Pan et al. (2024), who developed TAGExplainer, a method for generating natural language explanations for Text-Attributed Graphs (TAGs)[34]. Although their explanations incorporate elements of narrative communication, their framework is limited only to TAGs. This limitation restricts the generalisability of TAGExplainer to broader graph data modelling contexts, as TAGs represent only a subset of real-world graph data (see Appendix B.4 for examples).
@@ 3 Methods
3.1 XAI Narrative and Description
Research in psychology and communication theory indicates that narrative-based explanations are more accessible and memorable than descriptive methods, making them effective for conveying scientific evidence to non-expert audiences [6, 7, 3]. Moreover, narratives are processed more rapidly by individuals without prior knowledge and are more engaging and persuasive, thereby enhancing trust and understanding of AI models [31, 26, 3].
Further, as narrative communication relies on contextual cause-and-effect relationships, it is considerably more challenging to fragment a narrative into smaller, meaningful segments without either significantly altering the interpretation of these segments or disrupting the coherence of the original narrative [6]. Consequently, narratives are often perceived as storytelling, characterised by a coherent structure comprising an introduction, main body, and conclusion. In contrast, descriptive, context-free communication can be readily fragmented into smaller units while still effectively conveying the necessary information [6]. The XAIstories method addresses aforementioned explanation limitations by enhancing the narrative communication of SHAP and CF explanations, aligning with the research on human-AI interactions [26, 31, 3].
Having identified the need to distinguish between narrative and descriptive explanations in XAI, and drawing on social science research [11, 6, 31, 7], we propose definitions for both terms:
@@ 3.1 XAI Narrative and Description
In the context of explaining GNN models, a description would just list the most important features and neighbouring nodes. This relates closely to the data-to-text or graph-to-text approaches[35, 8]. Figure 4 presents an example of XAI Description for the subgraph and feature importance output provided by GNNExplainer shown in Figure 1. Clearly, in our context, a XAI Description is less valuable than an XAI Narrative, as descriptions are less accessible and memorable than narrative communication methods, making them less effective for conveying scientific evidence to a broader audience [6, 26].
3.2 From GNNs to Natural Language Narratives
To address the issue of purely technical explanations for GNN predictions, we propose GraphXAINs, natural language narratives for graphs. We propose the following definition of GraphXAIN, a XAI Narratives for graphs:
@@ 3.2 From GNNs to Natural Language Narratives
Our solution involves converting subgraph structures and feature importance metrics derived from graph explainers into GraphXAINs - coherent natural language narratives. The workflow, presented in Figure 2, begins with the training of a GNN model on a relational graph dataset for classification or regression tasks. Our approach is not limited to pure graphs; data can include associated node and edge features. Further, using the graph explainer method, we extract relevant subgraphs and feature importances. The subgraph, along with feature importance, are then transformed and prompted into an LLM for further inference. The LLM is guided to generate a natural language narrative that explains the prediction of the target node’s label or score, highlighting the contributions of specific neighbouring nodes, corresponding edges, and node features in a context-dependent manner, resulting in a coherent cause-and-effect narrative presented to the end user.
Importantly, our framework is agnostic to the graph data type, graph model, performed task (classification and regression), and graph explainer, allowing its application across various graph scenarios and applications. By translating technical explanations into coherent narratives, we aim to make graph model explanations more intuitive and accessible for practitioners.
@@ 4 Experiments
4.1 Datasets
We conduct our experiments on two real-world graph datasets, one used for the node classification task and the other for the node regression task. The dataset for the node classification task, NBA players’ dataset111https://www.kaggle.com/noahgift/social-power-nba, includes players’ performance statistics from one NBA season, alongside various personal attributes such as height, weight, age, and nationality. The graph is constructed by linking NBA players based on their relationships and interactions on the Twitter platform. After preprocessing, filtering out disconnected nodes, and ensuring the graph is undirected, the final graph comprised of the total number of 400 nodes (players), 42 node features, and 21,242 edges (Twitter connections). The node classification task involves predicting whether a player’s salary exceeds the median salary.
The dataset used to perform node regression task, IMDB movie dataset222https://www.kaggle.com/datasets/harshitshankhdhar/imdb-dataset-of-top-1000-movies-and-tv-shows, includes information on 1,000 movies from the IMDB database. The dataset includes information such as movie director, main actors, release year, duration, and genre. The graph is created by connecting nodes, represented by a single movie, by edges which link two movies if at least one actor played in both movies. In total, the graph consists of 1,000 nodes and 5,608 undirected edges. Each node has 12 features. The node regression task involves predicting the IMDB rating score ranging between 0 and 10, where 10 is the maximum score indicating the general appreciation of the movie by the audience.
4.2 Graph Models
Prior to training, both datasets were randomly divided into separate training, validation, and test sets in a 60/20/20 split. The GNN classification model consists of two Graph Convolutional Network (GCN) layers [21], with 16 hidden channels. For training, we use the Binary Cross-Entropy loss function with the AdamW [23] optimizer with a learning rate of 0.001 and a weight decay of 510-4. The training process continued for 1,400 epochs, resulting in the GCN classification model that achieved a test AUC of 0.80.
The GNN regressor model consists of two GCN layers with 32 hidden channels. The Root Mean Square Error loss function is used to train the model with the AdamW optimizer with a learning rate of 0.01 and a weight decay of 510-4. The training process continues for 7,500 epochs, with the early stopping of 500 steps, resulting in the RMSE of 0.28 on the held-out test set.